# Taylor polynomials on left-quotients of Carnot groups

### v1

@@ 1. Introduction
In the generality considered by Rothschild and Stein, one has to lift the vector fields to a free nilpotent Lie algebra, which is essentially the largest once the number of generators and the step are given. However, it is clear that in explicit examples, one can try to minimise the dimension of the manifold in which the lifted vector fields live. See, for example, [3].
In this article we reverse the logic. We start with vector fields in a stratified nilpotent Lie group and we project them onto sub-Riemannian manifolds. The crucial property of the projections that we consider is that they are submetries: metric balls of a fixed radius are mapped onto metric balls of the same radius. This allows to transfer several analytic properties of stratified Lie groups to possibly non-equiregular sub-Riemannian manifolds, where the span of horizontal vector fields bracket-generating the tangent space may have different dimensions at different points. A similar setup has been recently considered by [2].
@@
Therefore, we will transfer two more results that are linked to the Taylor polynomial theory: a sufficient condition for a function to be real analytic and the $L$-harmonicity of Taylor polynomials. These results will follow from the analogous results for stratified groups in [1]. We will do this in Section 5.
@@ 2. Preliminaries
We review the definitions of Carnot groups. We first consider stratified Lie groups and their Lie algebras from an algebraic point of view, and then equip a stratified Lie group with the sub-Riemannian (or Carnot–Carathéodory) distance to obtain a Carnot group. We then give the definition os sub-Riemannian manifold.
@@ 2.1. Carnot groups
We denote by $L_{g}$ the left translation by $g$ in $G$, that is, $L_{g}h=gh$ for all $h\in G$. Each $v$ in $\mathfrak{g}$ then induces a left-invariant vector field equal to $(L_{g})_{*}(v)$ at each point $g\in G$. The set $\overleftarrow{{\mathfrak{g}}}$ of all left-invariant vector fields with vector field commutation is isomorphic to $\mathfrak{g}$, and it inherits the stratification of $\mathfrak{g}$. The subbundle $HG$ of the tangent bundle $TG$, where $H_{g}G=(L_{g})_{*}(\mathfrak{g}_{1})$, is called the horizontal distribution. We will denote left invariant vector fields on $G$ by $\tilde{X}$. If $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$, then $\tilde{X}$ is homogeneous of degree $j$, in the sense that $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}\tilde{X}\,F\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function.
@@
where the infimum is taken over all horizontal curves $\gamma:[0,1]\to G$ such that $\gamma(0)=g$ and $\gamma(1)=g^{\prime}$. The stratified group $G$, equipped with the distance $d_{G}$, is known as a Carnot group. It is clear from the definition that left translations are isometries for $d_{G}$. Moreover, $d_{G}(\delta_{\lambda}(g),\delta_{\lambda}(g^{\prime}))=\lambda\,d_{G}(g,g^{\prime})$.
@@ 3. Right cosets and submetries
is a diffeomorphism from the Lie algebra onto $G$. We identify the points in $G$ using these exponential coordinates of the second kind. The identity of $G$ will then be the zero vector, denoted by 0. We then identify the quotient with $M=\mathbb{R}^{m}$ by
[…]
defined by $\iota(y_{1},\dots,y_{\ell},x_{1},\dots,x_{m})=(x_{1},\dots,x_{m})$. We abuse of the notation 0 for the point in $M$ such that $x_{j}=0$ for every $j$. The context will distinguish it from the identity of the Lie group $G$. Define a distance function on $M$ by $d_{M}=d\circ\iota^{-1}$. Let $\Pi=\iota\circ\pi$ and define its right inverse $\Phi^{-1}:M\to G$ so that $\Phi^{-1}(x_{1},\dots,x_{m})=(0,\dots,0,x_{1},\dots,x_{m})$. Clearly $\Pi$ is a submetry. In particular, the following holds.
@@
We will denote the left invariant vector fields in the coordinates introduced in (2) corresponding to the basis vectors $w_{1},\dots,w_{\ell}$ by $\tilde{Y}_{1},\dots,\tilde{Y}_{\ell}$ and those corresponding to the basis vectors $v_{1},\dots,v_{m}$ by $\tilde{X}_{1},\dots,\tilde{X}_{m}$. Let $f:M\to\mathbb{R}$ be a smooth function. The push-forward $X$ of a vector field $\tilde{X}$ on $G$ is defined by
@@ 3.1. Examples
Let $G$ be the connected and simply connected Lie groups with $\mathfrak{g}$ as Lie algebra. We use coordinates on $G$ such that
@@
which define a Grushin structure of order $\ell$ on the plane. If $p,q\in M$, and $g_{0}=\phi^{-1}(q)$, then Proposition 3.1 implies that there is $g\in G$ with $\Pi(g)=p$ so that $d_{M}(p,q)=d_{G}(g,g_{0})$. Moreover, if we define on $M$ the map $\delta_{\lambda}$ by
@@
with $\Pi:G\to M$ the usual projection. The distribution of vector fields $X_{1}$ and $X_{2}$ is a subbundle that defines a CR structure on $M$. Indeed, the latter can be identified with
@@ 4.1. Taylor polynomials on Carnot groups
Let $(G,d_{G})$ be a Carnot group and $\mathfrak{g}$ be its Lie algebra. We already observed that $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$ is homogeneous of degree $j$, that is $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}(\tilde{X}\,F)\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function. From the basis that we fixed in Section 3 we select $r$ vector fields that form a basis of left invariant vector fields of the first layer. It will be convenient to denote these by $\tilde{X}_{1},\dots,\tilde{X}_{r}$. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $\tilde{X}^{I}$ the differential operator defined as the product of $K$ left invariant vector fields each of homogeneous degree $d(i_{j})$. Then $\tilde{X}^{I}$ is homogeneous of degree $d(I)=d_{i_{1}}+\dots+d_{i_{K}}$. We write $\tilde{X}_{hor}^{I}$ if $d({i_{j}})=1$ for every $j$. We say that a function $F:G\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $F\in C^{k}_{hor}(G)$, if $\tilde{X}_{hor}^{I}\,F$ is continuous on $G$ for every $I$ with $d(I)\leq k$.
@@
Let $F:G\to\mathbb{R}$ be a $C_{hor}^{k}(G)$ function. The McLaurin polynomial of homogeneous degree $k$ related to $F$ is the unique polynomial $\tilde{P}=\tilde{P}_{k}(F,0)$ such that $\tilde{X}^{I}P(0)=\tilde{X}^{I}f(0)$, with $d(I)\leq k$. For $g_{0}\in G$ and $F\in C_{hor}^{k}(G)$, the polynomial
@@
There exist positive constants $c_{1}$ and $b$, depending only on the group $G$ and the distance $d_{G}$, such that for every $F\in C^{1}_{hor}(G)$,
$|F(gg_{0})-F(g)|\leq c_{1}d_{G}(g_{0},0)\sup\left\{|\tilde{X}_{j}F(gu)|\,:\,1\leq j\leq r,d_{G}(u,0)\leq b\,d_{G}(g_{0},0)\right\}$
@@
For every positive integer $k$ there is a constant $C_{k}$ such that for all $F\in C^{k}_{hor}(G)$ and all $g,g_{0}\in G$,
$|F(gg_{0})-\tilde{P}_{k}(F,g)(gg_{0})|\leq C_{k}d_{G}(g_{0},0)^{k}\eta(g,b^{k}d_{G}(g_{0},0)),$
@@
$\eta(g,b^{k}d_{G}(g_{0},0))=\sup\left\{|\tilde{X}_{hor}^{I}f(gu)-\tilde{X}_{hor}^{I}f(g)|\,:\,d(I)=k,\,d_{G}(u,0)\leq b^{k}d_{G}(g_{0},0)\right\}.$
@@
With the notation of Theorom 4.2 and $F\in C^{k+1}_{hor}(G)$,
$|F(gg_{0})-\tilde{P}_{k}(f,g)(gg_{0})|\leq C^{\prime}_{k}d_{G}(g_{0},0)^{k+1}\sup\left\{|\tilde{X}_{hor}^{I}f(gu)|\,:\,d(I)=k+1,\,d_{G}(u,0)\leq b^{k+1}d_{G}(g_{0},0)\right\},$
@@
4.2. Taylor polynomials for quotients and theorems
Let $(M,d_{M})$ be a sub-Riemannian manifold obtained as $\Pi(G)$ for some Carnot group $G$ and subgroup $H$, as defined in Section 3. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $X^{I}$ the differential operator defined as the product of $K$ images by $\Pi_{*}$ of left invariant vector fields on $G$, each of homogeneous degree $d(i_{j})$ (note that some of these operators may be zero). We write $X_{hor}^{I}$ if the vector fields involved are images through $\Pi_{*}$ of vector fields of homogeneous degree $1$. We say that a function $f:M\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $f\in C^{k}_{hor}(M)$, if $X_{hor}^{I}\,f$ is continuous on $M$ for every $I$ with $d(I)\leq k$. We will write $X_{j}=\Pi_{*}(\tilde{X}_{j}),j=1,\dots,r$.
Definition 4.2.
Let $f:M\to\mathbb{R}$ be a $C_{hor}^{k}(M)$ function. The McLaurin polynomial of homogeneous degree $k$ related to $f$ is the unique polynomial $P=P_{k}(f,0)$ defined by
$P_{k}(f,0)=\tilde{P}_{k}(f\circ\Pi,0).$
For $p\in M$ and $f\in C_{hor}^{k}(M)$, the polynomial
$P_{k}(f,p):=\tilde{P}_{k}(f\circ\Pi\circ L_{\Phi^{-1}(p)},0)\circ L^{-1}_{\Phi^{-1}(p)}$
is the Taylor polynomial of homogeneous degree $k$ related to $f$ around $p$.
Theorem 4.3 (Lagrange mean value theorem for quotients).
Given $f\in C^{1}_{hor}(M)$, there exist positive constants $c_{1}$ and $b$, depending only on the groups $G$ and $H$, and on the distance $d_{M}$, such that
(4) $|f(p)-f(q)|\leq c_{1}d_{M}(p,q)\sup\left\{|X_{j}f(s)|\,:\,1\leq j\leq r,d_{M}(s,q)\leq b\,d_{M}(p,q)\right\}$
@@ 4.2. Taylor polynomials for quotients and theorems
Let $F=f\circ\Pi$ and $g_{0}=\Phi^{-1}(q)$. Since $\Pi$ is a submetry, by Proposition 3.1 there are $g,u\in G$ with $\Pi(g)=p$ and $\Pi(u)=s$ so that $d_{G}(g,g_{0})=d_{M}(p,q)$ and $d_{G}(u,g_{0})=d_{M}(s,q)$. Formula (4) is equivalent to
$|f(\Pi(g))-f(\Pi(g_{0}))|\leq c_{1}d_{G}(g,g_{0})\sup\left\{|X_{j}f(\Pi(u))|\,:\,1\leq j\leq r,d_{G}(u,g_{0})\leq b\,d_{G}(g,g_{0})\right\}.$
Since $X_{j}f(\Pi(u))=X_{j}f(\Pi(u^{\prime}))$ if $\Pi(u)=\Pi(u^{\prime})$, the formula above is equivalent to
$|F(g)-F(g_{0})|\leq c_{1}d_{G}(g,g_{0})\sup\left\{|\tilde{X}_{j}F(u)|\,:\,1\leq j\leq r,d_{G}(u,g_{0})\leq b\,d_{G}(g,g_{0})\right\},$
which is the same as the conclusion of Theorem 4.1. Now we observe that $F\in C^{1}_{hor}(G)$ and the proof is finished. ∎
Theorem 4.4.
For every positive integer $k$ there is a constant $C_{k}$ such that for all $f\in C^{k}_{hor}(M)$ and all $p,q\in M$,
$|f(p)-P_{k}(f,q)(p)|\leq C_{k}d_{M}(p,q)^{k}\eta(q,b^{k}d_{M}(p,q)),$
where $b$ is as in the previous theorem and
$\eta(q,b^{k}d_{M}(p,q))=\sup\{|X_{hor}^{I}f(n)-X_{hor}^{I}f(q)|\mid d(I)=k,\,d_{M}(n,q)\leq b^{k}d_{M}(p,q)\}.$

### v3

@@ 1. Introduction
In the generality considered by Rothschild and Stein, one has to lift the vector fields to a manifold with symbol a free nilpotent Lie algebra, which is essentially the largest once the number of generators and the step are given. However, it is clear that in explicit examples, one can try to minimise the dimension of the manifold in which the lifted vector fields live. See, for example, [4], where a specific lifting is proposed and Taylor polynomial approximations are defined to define Hardy spaces in some non-equiregular sub-Riemannian manifolds. In this article we aim to generalise and clarify the method used in [4] (and other articles in preparation by the author and his collaborators) to obtain Taylor approximation theorems. We start with vector fields in a stratified nilpotent Lie group and we project them onto sub-Riemannian manifolds. The crucial property of the projections that we consider is that they are submetries: metric balls of a fixed radius are mapped onto metric balls of the same radius. This allows to transfer several analytic properties of stratified Lie groups to possibly non-equiregular sub-Riemannian manifolds, where the span of the horizontal vector fields bracket-generating the tangent space may have different dimensions at different points, or the growth vector may be different at different points. A similar setup has been recently considered by [3].
@@
Therefore, we will transfer two more results that are linked to the Taylor polynomial theory: a sufficient condition for a function to be real analytic and the $L$-harmonicity of Taylor polynomials of $L$-harmonic functions. These results will follow from the analogous results for stratified groups in [2]. We will do this in Section 5.
@@ 2. Preliminaries
We review the definitions of Carnot groups. We first consider stratified Lie groups and their Lie algebras from an algebraic point of view, and then equip a stratified Lie group with the sub-Riemannian (or Carnot–Carathéodory) distance to obtain a Carnot group. We then give the definition of sub-Riemannian manifold.
@@ 2.1. Carnot groups
We denote by $L_{g}$ the left translation by $g$ in $G$, that is, $L_{g}h=gh$ for all $h\in G$. Each $v$ in $\mathfrak{g}$ then induces a left-invariant vector field equal to $(L_{g})_{*}(v)$ at each point $g\in G$. The set $\overleftarrow{{\mathfrak{g}}}$ of all left-invariant vector fields with vector field commutations is isomorphic to $\mathfrak{g}$, and it inherits the stratification of $\mathfrak{g}$. The subbundle $HG$ of the tangent bundle $TG$, where $H_{g}G=(L_{g})_{*}(\mathfrak{g}_{1})$, is called the horizontal distribution. We will denote left invariant vector fields on $G$ by $\tilde{X}$. If $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$, then $\tilde{X}$ is homogeneous of degree $j$, in the sense that $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}\tilde{X}\,F\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function.
@@
where the infimum is taken over all horizontal curves $\gamma:[0,1]\to G$ such that $\gamma(0)=g$ and $\gamma(1)=g^{\prime}$. The stratified group $G$, equipped with the distance $d_{G}$, is called a Carnot group. It is clear from the definition that left translations are isometries for $d_{G}$. Moreover, $d_{G}(\delta_{\lambda}(g),\delta_{\lambda}(g^{\prime}))=\lambda\,d_{G}(g,g^{\prime})$.
@@ 3. Right cosets and submetries
is a diffeomorphism from the Lie algebra onto $G$. We identify the points in $G$ using these exponential coordinates of the second kind. The identity of $G$ will then be the zero vector, denoted by 0. We then consider the slice $M=\{g\in G\,:\,y_{1}=\dots=y_{\ell}=0\}\simeq\mathbb{R}^{m}$. The manifold $M$ is a global smooth cross-section for the projection, and therefore $\G\pi\big|_{M}:M\to H\backslash$ G is a diffeomorphism. Define a distance function on $M$ by $d_{M}=d\circ\pi$. Note that this distance is not equivalent to the distance induced by restricting the Carnot distance of $G$.
Let $\Pi=(\pi\big|_{M})^{-1}\circ\pi:G\to M$. Clearly $\Pi$ is a submetry. We take as its right inverse the inclusion mapping $(0,\dots,0,x_{1},\dots,x_{m})$ to itself, denoted $\iota$. In particular, the following holds (see [7]).
@@
We will denote the left invariant vector fields in the coordinates introduced in (2) corresponding to the basis vectors $w_{1},\dots,w_{\ell}$ by $\tilde{Y}_{1},\dots,\tilde{Y}_{\ell}$ and those corresponding to the basis vectors $v_{1},\dots,v_{m}$ by $\tilde{X}_{1},\dots,\tilde{X}_{m}$. A vector field $X$ on $M$ is $\pi$-related to a vector field $\tilde{X}$ on $G$ if $d\pi(\tilde{X})=d\pi(X)$. If $f:M\to\mathbb{R}$ is a smooth function, then
@@ 3.1. Examples
This is a filiform Lie algebra of step $\ell+1$. Let $G$ be the connected and simply connected Lie groups with $\mathfrak{g}$ as Lie algebra. We use coordinates on $G$ such that
@@
which define a Grushin structure of order $\ell$ on the plane. If $p,q\in M$, and $g_{0}=\iota(q)$, then Proposition 3.1 implies that there is $g\in G$ with $\Pi(g)=p$ so that $d_{M}(p,q)=d_{G}(g,g_{0})$. Moreover, if we define on $M$ the map $\delta_{\lambda}$ by
@@
with $\Pi:G\to M$ the usual projection. The distribution defined by the vector fields $X_{1}$ and $X_{2}$ is a subbundle that defines a CR structure on $M$. Indeed, the latter can be identified with
@@ 4.1. Taylor polynomials on Carnot groups
Let $(G,d_{G})$ be a Carnot group and $\mathfrak{g}$ be its Lie algebra. We already observed that $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$ is homogeneous of degree $j$, that is $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}(\tilde{X}\,F)\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function. From the basis that we fixed in Section 3 we select $r$ vector fields that form a basis of left invariant vector fields of the first layer. It will be convenient to denote these by $\tilde{X}_{1},\dots,\tilde{X}_{r}$. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $\tilde{X}^{I}$ the differential operator defined as the product of $K$ left invariant vector fields each of homogeneous degree $d(i_{j})$. Then $\tilde{X}^{I}$ is homogeneous of degree $d(I)=d_{i_{1}}+\dots+d_{i_{K}}$. We write $\tilde{X}_{\mathrm{hor}}^{I}$ if $d({i_{j}})=1$ for every $j$. We say that a function $F:G\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $F\in C^{k}_{\mathrm{hor}}(G)$, if $\tilde{X}_{\mathrm{hor}}^{I}\,F$ is continuous on $G$ for every $I$ with $d(I)\leq k$.
@@
Let $F:G\to\mathbb{R}$ be a $C_{\mathrm{hor}}^{k}(G)$ function. The McLaurin polynomial of homogeneous degree $k$ related to $F$ is the unique polynomial $\tilde{P}=\tilde{P}_{k}(F,0)$ such that $\tilde{X}^{I}P(0)=\tilde{X}^{I}f(0)$, with $d(I)\leq k$. For $g_{0}\in G$ and $F\in C_{\mathrm{hor}}^{k}(G)$, the polynomial
@@
There exist positive constants $c_{1}$ and $b$, depending only on the group $G$ and the distance $d_{G}$, such that for every $F\in C^{1}_{\mathrm{hor}}(G)$,
$|F(gg_{0})-F(g)|\leq c_{1}d_{G}(g_{0},0)\sup\left\{|\tilde{X}_{j}F(gu)|\,:\,1\leq j\leq r,d_{G}(u,0)\leq b\,d_{G}(g_{0},0)\right\}.$
@@
For every positive integer $k$ there is a constant $C_{k}$ such that for all $F\in C^{k}_{\mathrm{hor}}(G)$ and all $g,g_{0}\in G$,
$|F(gg_{0})-\tilde{P}_{k}(F,g)(gg_{0})|\leq C_{k}d_{G}(g_{0},0)^{k}\tilde{\eta}_{G}(g,b^{k}d_{G}(g_{0},0)),$
@@
$\tilde{\eta}_{G}(g,b^{k}d_{G}(g_{0},0))=\sup\left\{|\tilde{X}_{\mathrm{hor}}^{I}F(gu)-\tilde{X}_{\mathrm{hor}}^{I}F(g)|\,:\,d(I)=k,\,d_{G}(u,0)\leq b^{k}d_{G}(g_{0},0)\right\}.$
@@
With the notation of Theorom 4.2 and $F\in C^{k+1}_{\mathrm{hor}}(G)$, there is $C_{k}^{\prime}$ so that
$|F(gg_{0})-\tilde{P}_{k}(F,g)(gg_{0})|\leq C^{\prime}_{k}d_{G}(g_{0},0)^{k+1}\sup\left\{|\tilde{X}_{\mathrm{hor}}^{I}F(gu)|\,:\,d(I)=k+1,\,d_{G}(u,0)\leq b^{k+1}d_{G}(g_{0},0)\right\}.$
@@
4.2. Taylor polynomials for quotients and theorems.
Let $(M,d_{M})$ be a sub-Riemannian manifold obtained as the image $\Pi(G)$ of a Carnot group $G$ and subgroup $H$, as defined in Section 3. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $X^{I}$ the differential operator defined as the product of $K$ images by $\Pi_{*}$ of left invariant vector fields on $G$, each of homogeneous degree $d(i_{j})$ (note that some of these operators may be zero). We write $X_{\mathrm{hor}}^{I}$ if the vector fields involved are images through $\Pi_{*}$ of vector fields of homogeneous degree $1$. We say that a function $f:M\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $f\in C^{k}_{\mathrm{hor}}(M)$, if $X_{\mathrm{hor}}^{I}\,f$ is continuous on $M$ for every $I$ with $d(I)\leq k$. A polynomial of degree $k$ is a function $P:M\to\mathbb{R}$ so that $X_{\mathrm{hor}}^{I}P$ is not identically zero for some $I$ with $d(I)=k$ and $X_{\mathrm{hor}}^{I}P=0$ for all multi-index such that $d(I)\geq k+1$.
[…]
Definition 4.2 (Taylor polynomials on $M$).
Let $f\in C_{\mathrm{hor}}^{k}(M)$ and $q\in M$. The Taylor polynomial of degree $k$ of $f$ at $q$, denoted $P_{k}(f,q)$, is the unique polynomial $P:M\to\mathbb{R}$ such that
$X_{\mathrm{hor}}^{I}P(q)=X_{\mathrm{hor}}^{I}f(q)\qquad\text{for every multi-index }I\text{ with }d(I)\leq k.$
If $q=0$, we will call $P$ the McLaurin polynomial.
The following Lemma shows that the definition of Taylor polynomial on $M$ is well posed.
Lemma 4.1.
Let $Q$ be a polynomial on $M$ of degree $\leq k$. If $X_{\mathrm{hor}}^{I}Q(q)=0$ for all multi-indices $I$ with $d(I)\leq k,$ then $Q\equiv 0$ on $M$.
@@
Let $\tilde{Q}:=Q\circ\Pi$. Then $\tilde{Q}$ is a left HH–invariant homogeneous polynomial on $G$. For any multi-index $I$ and any representative $g_{0}$ of $q$, we have
$\tilde{X}^{I}\tilde{Q}(g_{0})=X^{I}Q(q).$
By assumption, $\tilde{X}_{\mathrm{hor}}^{I}\tilde{Q}(g_{0})=0$ for all $|I|\leq k$. By the uniqueness of Taylor polynomials on $G$, this implies $\tilde{Q}\equiv 0$. Restricting to $M$ yields $Q\equiv 0$ on $M$. ∎
In the next proposition we relate the Taylor polynomials on $M$ to those on $G$. This will be crucial for proving remainder estimates.
Proposition 4.1.
Let $f\in C_{\mathrm{hor}}^{k}(M)$ and $\tilde{f}:=f\circ\Pi$. Fix $q\in M$ and choose any $g_{0}\in G$ with $\Pi(g_{0})=q$. Let $\tilde{P}_{k}(\tilde{f},g_{0})$ denote the Taylor polynomial of homogeneous degree $k$ for $\tilde{f}$ at $g_{0}$. Then:
  1. (1)
[…]
  2. (2)
    The restriction of $\bar{P}_{k}$ to $M$ (via the identification $\psi$) coincides with the Taylor polynomial on $M$, namely
    $P_{k}(f,q)\;=\;\bar{P}_{k}\circ\psi\;=\;\tilde{P}_{k}(\tilde{f},g_{0})\big|_{M}.$
  3. (3)
    The polynomial $\tilde{P}_{k}(\tilde{f},g_{0})$ (and hence $\bar{P}_{k}$ and $P_{k}(f,q)$) is independent of the choice of representative $g_{0}$ with $\Pi(g_{0})=q$.